# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v2

@@ 1 Minimalist Market Design: A Proven Framework
The increasing prominence of market design in mainstream economics has been recognized through several notable awards. In 2012, Alvin Roth and Lloyd Shapley shared the Nobel Memorial Prize in Economic Sciences for their contributions to “Stable Allocations and the Practice of Market Design.” Parag Pathak received the 2018 John Bates Clark Medal for his work in market design and the economics of education. In 2020, Paul Milgrom and Robert Wilson shared the Nobel Memorial Prize in Economic Sciences for their Improvements to “Auction Theory and Inventions of New Auction Formats.” As Jackson (2013) notes, “this is an area where microeconomic theory has had its largest direct impact.”
@@ 1.1 What is Minimalist Market Design?
Similarly, in minimalist market design, a design economist starts by identifying the mission of the institution—the primary objectives of decision makers, system operators, or other stakeholders involved in the institution’s design. Understanding the institution’s historical evolution is often crucial for recognizing these objectives.
Next, the economist assesses whether the current setup meets these objectives and fulfills the institution’s mission. If not, this misalignment signals an opportunity for policy improvement through better design.
@@
In the late 1990s, as a junior design economist, I observed that emerging market design approaches—for example, the engineering approach to market design (Roth and Peranson, 1999; Roth, 2002)—predominantly, if not exclusively, reflected the experiences of leading economists commissioned to design or reform institutions. [^2] Despite significant advancements in the field, this trend still persists. The commissioning of such projects by decision-makers indicates their readiness for design or reform, often driven by urgent needs or crises, which makes them particularly receptive to proposed changes.
@@
Although Medawar’s concern focuses on the discovery process in scholarly inquiry, similar issues arise when attributing policy successes to specific scholarly research in applied economics, including market design. Often, the policy impact attributed to market design scholarship reflects ”post hoc” scholarly progress occurring after the practical impact, especially when the design team is commissioned for the project. In contrast, throughout this monograph, I will be mindful of Medawar’s concerns by presenting the thought processes that accompany minimalist market design. I will not only focus on applications where policy impact has occurred following the initial research findings but also present research and policy in a holistic manner that captures their interactions and mutual evolution.
@@
Sometimes, practical impact follows the research but occurs independently of it, initiated by policymakers or other stakeholders who discover and pursue similar solutions. In the minimalist market design framework—where a key motivation is uncovering the institution’s intended design—this scenario is particularly likely. Strictly speaking, I do not classify such outcomes as policy impact. Instead, if decision makers independently act in ways consistent with the minimalist market design previously outlined in scholarly work—arriving at similar conclusions and implementing changes—I refer to this as external validity. So, whenever I speak of external validity, “life imitates science.”
While achieving direct policy impact may be considered by many as the more “glorifying” outcome, I personally believe that external validity is at least as important and serves a unique role. Achieving policy impact—though much less likely for an outsider critic than for a commissioned insider—can be influenced by factors unrelated to the merits of the proposed design, such as personal charisma, strong connections, an urgent need to avert a crisis, or sheer luck. In contrast, external validity serves as direct proof that the minimalist framework is working exactly as intended. It strongly confirms that you have indeed addressed the root causes of a flawed institution’s failure and formulated a desirable improvement.
@@ 1.3 Evolution of Minimalist Market Design
Phase 3—currently ongoing—represents my deliberate and systematic application of minimalist market design to shape both research and policy efforts, beginning in the early 2010s. During this phase, I have applied the framework across various areas, often successfully, though not always, in collaboration with diverse teams, including:
@@ 1.4 Aims, Theses and Outline of the Monograph
First, I aim to outline a coherent institutional design framework—the minimalist market design—which adopts a holistic approach to research and policy. Intended for policy-oriented scholars who aspire to influence policy from outside the system, this framework enables even outsider critics to build persuasive cases for change grounded in practical, use-inspired basic research. The thesis associated with this aim is that by adopting minimalist market design, these scholars can increase the likelihood of their proposals being adopted and implemented by decision-makers.
@@
Third, I aim to provide a detailed intellectual history of key innovations and policy breakthroughs in matching markets in the last 25 years. By highlighting the role of minimalist market design in these achievements, I present this history through a series of discovery-invention cycles, as described by Narayanamurti and Odumosu (2016), which connect these innovations.
@@ 2.1 Navigating the Challenges of Policy Influence as an Outsider Critic
Let’s begin by exploring the motivation behind minimalist market design. Why did I, as an aspiring design economist, choose to venture into uncharted territory by experimenting with an untested approach rather than following mainstream methods?
At first glance, this path might seem unconventional, but I had compelling reasons for adopting this strategy in my quest to influence policy through research. Producing work that drives meaningful policy changes has always been a higher priority for me than work aimed solely at an academic audience. This focus naturally led my approach to diverge from the methods favored by the field’s pioneers.
At that time, most market design literature and methodologies were derived from commissioned design or reform projects, where the need for change was already recognized and expert input was actively sought—often in response to an urgent crisis. As an aspiring design economist, I realized that such opportunities might not present themselves to me for a long time, if ever, posing a significant challenge to my policy ambitions. Moreover, I believed that these experiences were less informative for situations where decision makers had not yet acknowledged the need for change and were likely to have vested interests in preserving the status quo.
This is not to say that commissioned market design projects do not yield significant theoretical contributions, which were central to my focus. They certainly do, much like the valuable scholarly work produced from these projects using other methods. However, these contributions often come after the project’s completion, reflecting lessons learned during or after its execution. As a result, many of these contributions—finalized post-project—are directed more toward an academic audience than practitioners. More importantly, it was not this research that made the original commission possible.
Therefore, I questioned whether lessons from commissioned market design projects—typically written for an academic audience—would be equally effective for critics attempting to inform policy from outside the system. While they might be, I had my doubts. Consequently, I developed a framework tailored to the needs of an aspiring design economist offering unsolicited policy advice as an outsider, often challenging established institutions.
@@ 2.2 Three Pillars of Minimalist Market Design
    Avoid Nonessential Changes. As an outsider critic, you may not fully understand how every component within an institution functions. Therefore, it is prudent to avoid altering elements that do not impact the primary objectives, thus minimizing the risk of unintended consequences.
@@
Based on these pillars, a minimalist design economist addressing the failures of a flawed institution first identifies the key policy objectives, then determines the root causes of the failures, and finally addresses them by directly targeting their sources. To maximize practical relevance and convince stakeholders of the value of these ideas, minimalist design economists typically develop custom models tailored to specific applications and create novel theories to identify and resolve the root causes of failures.
@@ 2.2.3 Forging a Design Framework with a Built-in Persuasion Strategy
Consistent with the first two pillars of minimalist market design, once a realistic model is crafted, the proposed solutions must closely align with the aims of policymakers and other stakeholders without compromising any unidentified objectives. Essentially, it must be demonstrated that the proposed design better meets the stakeholders’ goals than the current system. Such reforms provide genuine value to decision makers and are easier to justify to superiors or constituents.
@@
Relatedly, except in very rare cases, minimalist market design almost always begins with a practical problem to solve; modeling and analysis follow. You cannot realistically introduce and solve an analytical model and then hope that it addresses a practical problem of significance. The motivating practical problem must be taken very seriously. In this sense, underlying policy aspirations shape the research questions in minimalist market design. This approach stands in sharp contrast to most theory papers, where the application is treated merely as an ex-post reflection or an excuse to write down a model the authors wanted to solve.
@@
2.3 Features of Minimalist Market Design
In light of these pillars, minimalist market design embodies several key features that operationalize its foundational principles:
@@ 2.3 Features of Minimalist Market Design
  • Value for Aspiring Market Designers: Its built-in persuasion strategy makes it particularly valuable for aspiring design economists in settings where stakeholders have not yet recognized the need for change.
@@
  • Axiomatic Methodology: Often employs the axiomatic methodology as the principal technique for formal analysis.
In essence, these features of minimalist market design operationalize its three foundational pillars, offering a practical framework for effective institutional reform.
@@ 3 From Theory to Practice: The Role of Innovation Models
This perspective largely reflects an engineering approach to market design, first discussed and advocated by Roth and Peranson (1999) and Roth (2002). Under this approach, theory—typically “off-the-shelf” and developed without specific consideration for practical use—plays a less prominent role than in minimalist market design, primarily offering intuition and guiding practical designs in the right direction. They describe this role as “less heroic,” a perspective frequently echoed by other field leaders involved in major commissioned market design projects:
@@
This perspective—favoring simple models over those tailored specifically to individual projects—is particularly compelling for commissioned work under tight deadlines, where developing novel theoretical contributions is often impractical. As Narayanamurti and Odumosu (2016) highlight:
@@
Indeed, Roth (2002)—in his seminal work describing the foundations and early successes of the engineering approach to market design—emphasizes:
“First, design is often required to be fast. In each of these three cases, only about a year elapsed between the commissioning of a new market design and its delivery.” (p. 1345)
In my view, the relatively indirect relationship between theory and practice in engineering approach to market design is largely shaped by the features and demands of commissioned projects. Design economists working on such projects lack the flexibility to choose their problems or timelines; they must deliver solutions that address the complexities of the project while meeting strict deadlines. As Narayanamurti and Odumosu (2016) highlight, the pressures of completing a “scheduled” activity often leave little room for advancing the boundaries of knowledge. This dynamic, however, is not unique to market design; it reflects a broader paradigm articulated by Bush (1945), which promotes a linear, unidirectional relationship between science and technological innovation—a model that has profoundly influenced U.S. science and technology policy since World War II.
@@
Consequently, without these constraints, minimalist market design emerged as an exception to the broader trend influenced by Bush’s linear model of science and technological innovation. It diverges by embracing a holistic approach to research and practice, aligning with alternative models of innovation proposed by Stokes (1997) and Narayanamurti and Odumosu (2016). To illustrate these differences, we will first review Bush’s post-war science and technology framework in Section 3.1, followed by Stokes’s model in Section 3.2, and the model proposed by Narayanamurti and Odumosu in Section 3.3.
@@ 3.2 Beyond the Linear Model: Pasteur’s Quadrant
Within this classification, much of the research in economic theory, including mechanism design and information design, falls into Bohr’s Quadrant. In contrast, research in the engineering approach to market design and artificial intelligence belongs to Edison’s Quadrant. Bush’s linear model of innovation is most closely associated with these two quadrants.
@@
To better understand how these quadrants manifest in the field of design economics, let us explore specific examples that illustrate research practices within each quadrant. We will focus on seminal works that have not only advanced design economics but have also played a role in their authors receiving Nobel Prizes, thereby underscoring their significance. We begin with Bohr’s Quadrant.
@@ 3.2.1 Bohr’s Quadrant: Nash Implementation
The power of this result lies in its generality. However, it is evident that this mechanism is not intended for practical application. Beyond the complexity of Maskin’s mechanism, the broader concept of implementation via a Hurwicz-style mechanism—relying on achieving equilibrium outcomes under various rationality assumptions—renders Maskin’s exercise fairly abstract. This places the exercise, like many others in the mechanism design literature, firmly within Bohr’s Quadrant.
While such theoretical contributions are invaluable for expanding fundamental understanding, they do not directly address practical problems. This contrasts with research situated in Edison’s Quadrant, which focuses on solving real-world issues using applied research methods. An exemplary case of this is the redesign of the U.S. Medical Residency Match.
@@
Amid mounting dissatisfaction and the threat of market failure, the NRMP commissioned Alvin Roth in the fall of 1995 to redesign the matching process. Roth collaborated with Elliott Peranson, President of NMS—the private contractor managing the NRMP’s medical match at the time—to develop a new heuristic algorithm known as the Roth and Peranson design. This design was based on Gale and Shapley’s celebrated individual-proposing DA algorithm for a simplified version of the problem but was adapted to address the complexities specific to the NRMP’s situation.
@@
With the new design, addressing the first concern raised by Williams (1995) was straightforward. Simply replacing the institution-proposing DA algorithm with the individual-proposing DA algorithm as the starting point of the matching process resolved this issue. To tackle potential instability arising from match variations such as dual-physician couples, Roth and Peranson adopted an engineering approach. They leveraged an earlier technique by Roth and Vande Vate (1990), which had proven effective in simpler settings. The expectation was that this approach would perform reasonably well even in the more complex current setting, given that the fraction of residents in the system who were members of dual-physician couples was relatively low.
@@
These findings confirmed the earlier suggestion by Peranson and Randlett (1995) that the criticisms in Williams (1995) were overblown and that the initial starting outcome has little impact in any matching process that reaches a stable outcome—a near certainty given their findings. Nevertheless, the new Roth and Peranson design chose the individual-proposing DA algorithm as its starting point to remove any doubt about the fairness of the procedure for the residents. Even though the new design might generate outcomes very similar to those of the previous system that raised concerns, by addressing these issues, the new design would be a success. After all, Roth was called in to help ease the confidence crisis in the process.
@@
Klaus and Klijn (2005) examined the preference structures of couples that guarantee the existence of a stable outcome. They identified responsive preferences as one such domain.This structure is particularly plausible in settings where programs are in close proximity to one another. It assumes that each member of a couple has their own preferences over individual programs, and their joint preferences are structured such that, when the assignment of one member is fixed, an improvement for the other member based on their individual preferences also improves the couple’s joint preferences. [^20] Responsive preferences were originally introduced by Roth (1985) in a two-sided matching market to derive institution preferences over groups of individuals from their preferences over individual members, in a class of settings without complementarities. In such settings, the individual-proposing DA algorithm, which disregards joint preferences and treats couples as separate individuals, generates a stable outcome.
@@
Subsequent literature illuminated the reasons behind the virtually guaranteed existence of stable outcomes in the computational experiments with simulated data reported in Roth and Peranson (1999). These experiments reflected the small proportion of residents in the NRMP pool who are members of dual-physician couples. Kojima et al. (2013) established that, with random preferences and a vanishingly small proportion of couples, a stable outcome almost always exists. However, in the same model, Ashlagi et al. (2014) showed that as the proportion of couples increases, the probability of a stable outcome existing declines significantly.
@@
These theoretical advances, inspired by practical challenges, illustrate how applied research can also contribute to fundamental understanding. This convergence of basic and applied research—use-inspired basic research—embodies the essence of Pasteur’s Quadrant. To further illustrate this, we turn to William Vickrey’s foundational work in auction theory.
3.2.3 Pasteur’s Quadrant: Revenue Equivalence Theorem and Vickey Auction

### v3

@@ 1 Minimalist Market Design: A Proven Framework
The increasing prominence of market design in mainstream economics has been recognized through several notable awards. In 2012, Alvin Roth and Lloyd Shapley shared the Nobel Memorial Prize in Economic Sciences for their contributions to “Stable Allocations and the Practice of Market Design.” Parag Pathak received the 2018 John Bates Clark Medal for his work in market design and the economics of education. In 2020, Paul Milgrom and Robert Wilson shared the Nobel Memorial Prize in Economic Sciences for their contributions to “Auction Theory and Inventions of New Auction Formats.” As Jackson (2013) notes, “this is an area where microeconomic theory has had its largest direct impact.”
@@ 1.1 What is Minimalist Market Design?
This is exactly how a design economist organizes their efforts and operates within the minimalist market design framework. [^2] Throughout this monograph, the term design economist is used interchangeably with market designer to refer to researchers in the market design community who use formal methods to study or design institutions. They are not necessarily trained as economists; many, in fact, come from fields such as computer science, operations research, and electrical engineering.
They start by identifying the mission of the institution—the primary objectives of decision makers, system operators, or other stakeholders involved in the institution’s design. Understanding the institution’s historical evolution is often crucial for recognizing these objectives.
Next, the economist assesses whether the current setup meets these objectives and fulfills the institution’s mission. If not, this misalignment presents an opportunity for policy improvement through better design.
@@
In the late 1990s, as a junior design economist, I observed that emerging market design approaches—for example, the engineering approach to market design (Roth and Peranson, 1999; Roth, 2002)—predominantly, if not exclusively, reflected the experiences of leading economists commissioned to design or reform institutions. [^3] Despite significant advancements in the field, this trend still persists. The commissioning of such projects by decision makers indicates their readiness for design or reform, often driven by urgent needs or crises, which makes them particularly receptive to proposed changes.
@@
Although Medawar’s concern focuses on the discovery process in scholarly inquiry, similar issues arise when attributing policy successes to specific scholarly research in applied economics, particularly in market design. Often, the policy impact attributed to market design scholarship reflects post hoc scholarly progress that follows practical outcomes, especially when the design team is commissioned for the project. In this monograph, I will remain mindful of Medawar’s concerns and present the thought processes underlying minimalist market design. To this end, I will focus on applications where policy impact follows the initial research findings, presenting research and policy as a holistic interplay that captures their interactions and mutual evolution.
@@
Sometimes, practical relevance follows the research but occurs independently of it, initiated by policymakers or other stakeholders who discover and pursue similar solutions. In the minimalist market design framework—where a key motivation is uncovering the institution’s intended design—this scenario is particularly likely. Strictly speaking, I do not classify such outcomes as policy impact. Instead, if decision makers independently act in ways consistent with the minimalist market design previously outlined in scholarly work—arriving at similar conclusions and implementing changes—I refer to this as external validity. So, whenever I speak of external validity, “life imitates science.”
While achieving direct policy impact may be considered by many as the more “glorifying” outcome, I personally believe that external validity is at least as important and serves a unique role. Achieving policy impact—though much less likely for an outsider critic than for a commissioned insider—can be influenced by factors unrelated to the merits of the proposed design, such as personal charisma, strong connections, an urgent need to avert a crisis, or sheer luck. In contrast, external validity serves as direct proof that the minimalist framework is working exactly as intended. It strongly confirms that you have indeed addressed the root causes of a flawed institution’s failures and formulated a desirable improvement.
@@ 1.3 Evolution of Minimalist Market Design
Phase 3—currently ongoing—marks my deliberate and systematic application of minimalist market design to shape both my research and policy initiatives, which began in the early 2010s. During this phase, I have applied the framework across various areas—often successfully, though not always—in collaboration with diverse teams, including:
@@ 1.4 Aims, Theses and Outline of the Monograph
First, I aim to outline a coherent institutional design framework—the minimalist market design—which adopts a holistic approach to research and policy. Intended for policy-oriented scholars who aspire to influence policy from outside the system, this framework enables even outsider critics to build persuasive cases for change grounded in practical, use-inspired basic research. The thesis associated with this aim is that by adopting minimalist market design, these scholars can increase the likelihood that their proposals will be adopted by decision-makers and implemented in practice.
@@
Third, I aim to provide a detailed intellectual history of key innovations and policy breakthroughs in the theory and practice of matching markets over the past 25 years. By connecting these innovations and breakthroughs through a series of Discovery-Invention Cycles à la Narayanamurti and Odumosu (2016), I also highlight the role of minimalist market design in these achievements.
@@ 2.1 Navigating the Challenges of Policy Influence as an Outsider Critic
Let’s begin by exploring the motivation behind minimalist market design. Why did I, as an aspiring design economist, venture into uncharted territory and experiment with an untested approach, rather than follow the methods established by the pioneers of this emerging field?
At first glance, this path might seem unconventional. However, my decision was driven by a clear priority: to produce research that drives meaningful policy change rather than work aimed solely at an academic audience. This focus naturally steered me toward a different approach, one that diverged from the methods favored by the field’s pioneers.
At that time, most market design literature and methodologies focused on commissioned design or reform projects, where the need for change was already recognized, and expert input was actively sought—often in response to an urgent crisis. As an aspiring design economist, I realized that such opportunities might not come my way for a long time, if ever, posing a significant challenge to my policy ambitions. Moreover, I believed these experiences were less informative for situations where decision-makers had yet to acknowledge the need for change and were likely to have vested interests in preserving the status quo.
This is not to say that commissioned market design projects do not yield significant theoretical contributions, which were central to my focus. They certainly do, much like the valuable scholarly work produced from these projects using other methods. However, these contributions often come after the project’s completion, reflecting lessons learned during or after its execution. As a result, many of these contributions—finalized post-project—are directed more toward an academic audience than practitioners. More importantly, it was not this research that enabled the original commission.
Therefore, I questioned whether lessons from commissioned market design projects—typically written for an academic audience—would be equally effective for critics like myself seeking to inform policy from outside the system. While they might be, I had my doubts. Consequently, I developed a framework tailored to the needs of an aspiring design economist offering unsolicited policy advice as an outsider, often challenging established institutions.
@@ 2.2 Three Pillars of Minimalist Market Design
    Avoid Nonessential Changes. As an outsider critic, you may not fully understand the role of every aspect of an institution. Therefore, it is prudent to avoid altering elements that do not affect the primary objectives, minimizing the risk of unintended consequences.
@@
Based on these pillars, minimalist design economists addressing the failures of a flawed institution first identify the key policy objectives, determine the root causes of the failures, and then address them by directly targeting their sources. To maximize practical relevance and persuade stakeholders of the value of these ideas, they typically develop custom models tailored to specific applications and create novel theories to identify and resolve the root causes of failure.
@@ 2.2.3 Forging a Design Framework with a Built-in Persuasion Strategy
Consistent with the first two pillars of minimalist market design, once a realistic model is crafted, the proposed solutions must closely align with the aims of policymakers and other stakeholders, ideally without inadvertently compromising any unidentified objectives. Essentially, it must be demonstrated that the proposed design better meets the stakeholders’ goals than the current system. Such reforms provide genuine value to decision makers and are easier to justify to superiors or constituents.
@@
Relatedly, except in very rare cases, minimalist market design almost always begins with a practical problem to solve; modeling and analysis come next. You cannot realistically introduce and solve an analytical model and then hope it happens to address a significant practical problem. The motivating practical problem must be taken very seriously. In this sense, underlying policy aspirations shape the research questions in minimalist market design. This approach stands in sharp contrast to most theory papers, where the application is treated merely as an ex-post reflection or an excuse to write down a model the authors wanted to solve.
@@
2.3 Characteristics and Features of Minimalist Market Design
In light of these pillars, minimalist market design embodies several characteristics that operationalize its foundational principles:
@@ 2.3 Characteristics and Features of Minimalist Market Design
Taken together, these characteristics operationalize the three foundational pillars of minimalist market design, providing a practical framework for effective institutional reform.
While not a universal element of the framework, many of the practical applications presented in this monograph utilize an additional methodological feature to capture the institutional mission:
  • Axiomatic Methodology: Often (though not always) employs the axiomatic methodology as a central technique for formal analysis.
@@ 3 From Theory to Practice: The Role of Innovation Models
This perspective largely reflects an engineering approach to market design, first discussed and advocated by Roth and Peranson (1999) and Roth (2002). Under this approach, theory—typically “off-the-shelf” and developed without specific consideration for practical use—plays a less prominent role than in minimalist market design, primarily offering intuition and guiding practical designs in the right direction. Roth and Wilson describe this role as “less heroic,” a perspective frequently echoed by other field leaders involved in major commissioned market design projects:
@@
This perspective—favoring simple models over those tailored specifically to individual projects—is particularly compelling for commissioned projects under tight deadlines, where developing novel theoretical contributions is often impractical. As Narayanamurti and Odumosu (2016) highlight:
@@
The engineering approach to market design is primarily associated with such commissioned projects. For example, Roth (2002), in his seminal work on the foundations and early successes of market design, emphasizes:
“Aside from technical issues, these design efforts also shared some features that seem to be characteristic of the context in which markets are designed. First, design is often required to be fast. In each of these three cases, only about a year elapsed between the commissioning of a new market design and its delivery.” (p. 1345)
The indirect relationship between theory and practice in the engineering approach to market design is largely shaped by the demands of commissioned projects. Design economists in such settings lack the flexibility to choose their problems or timelines; they must deliver practical solutions under strict deadlines. As Narayanamurti and Odumosu (2016) note, the pressures of completing a “scheduled” activity often leave little room for advancing theoretical boundaries. This dynamic, however, is not unique to engineering approach to market design—it reflects a broader paradigm articulated by Bush (1945), which envisions a linear, unidirectional relationship between scientific discovery and technological innovation. This model has profoundly influenced U.S. science and technology policy since World War II.
@@
Freed from these constraints, minimalist market design emerged as an exception to the broader trend shaped by Bush’s linear model. It embraces a holistic approach that integrates research and practice, aligning instead with alternative innovation frameworks advocated by Stokes (1997) and Narayanamurti and Odumosu (2016). To clarify these distinctions, we will first review Bush’s post-war science and technology framework in Section 3.1, then examine Stokes’s Pasteur’s Quadrant model in Section 3.2, and finally discuss the Discovery-Invention Cycles model introduced by Narayanamurti and Odumosu in Section 3.3.
@@ 3.2 Beyond the Linear Model: Pasteur’s Quadrant
Within this classification, much of the research in economic theory, including mechanism design and information design, falls into Bohr’s Quadrant. In contrast, research in the engineering approach to market design and artificial intelligence belongs to Edison’s Quadrant. Bush’s linear model of innovation is most closely associated with these two quadrants, though research in Edison’s Quadrant often includes product development as well.
@@
To better understand how these quadrants manifest in the field of design economics, let us explore specific examples that illustrate research practices within each quadrant. We will focus on seminal works that have not only advanced the field of design economics but have also contributed to their authors receiving the Nobel Memorial Prize in Economic Sciences, thereby underscoring their significance. We begin with Bohr’s Quadrant.
@@ 3.2.1 Bohr’s Quadrant: Nash Implementation
The power and elegance of this result lie in its generality. However, it is evident that this mechanism is not intended for practical use. Beyond the complexity of Maskin’s mechanism, the broader concept of implementation via a Hurwicz-style mechanism—relying on achieving equilibrium outcomes under various rationality assumptions—renders Maskin’s exercise fairly abstract. This places the exercise, like many others in the mechanism design literature, firmly within Bohr’s Quadrant.
While such theoretical contributions are invaluable for expanding fundamental understanding, they do not directly address practical problems. This contrasts with research situated in Edison’s Quadrant, which focuses on solving real-world issues through applied research methods. A representative example of this is the redesign of the U.S. Medical Residency Match.
@@
Amid mounting dissatisfaction and the threat of market failure, the NRMP commissioned Alvin Roth in the fall of 1995 to redesign the matching process. Roth collaborated with Elliott Peranson to develop a new heuristic algorithm. These efforts culminated in the Roth and Peranson design, a procedure based on Gale and Shapley’s celebrated individual-proposing DA algorithm for the simplified version of the problem, but modified to address the complexities specific to the NRMP’s situation.
@@
With the new design, addressing the first concern raised by Williams (1995) was straightforward. Simply replacing the institution-proposing DA algorithm with the individual-proposing DA algorithm as the starting point of the matching process would resolve this issue. To tackle potential instability arising from match variations such as dual-physician couples, Roth and Peranson adopted an engineering approach. They leveraged an earlier technique by Roth and Vande Vate (1990), which had proven effective in simpler settings. The expectation was that this approach would perform reasonably well even in the more complex current setting, given that the fraction of residents in the system who were members of dual-physician couples was relatively low.
@@
These findings confirmed the earlier suggestion by Peranson and Randlett (1995) that the criticisms in Williams (1995) were overblown and that the initial starting outcome has little impact in any matching process that reaches a stable outcome—a near certainty given their findings. Nevertheless, the new Roth and Peranson design chose the individual-proposing DA algorithm as its starting point to remove any doubt about the fairness of the procedure for the residents. Even though the new design might generate outcomes very similar to those of the previous system that raised concerns, by addressing these concerns, the new design would be a success. After all, Roth was called in to help ease the confidence crisis in the process.
@@
Klaus and Klijn (2005) examined the preference structures of couples that guarantee the existence of a stable outcome. They identified responsive preferences as one such domain. This structure is particularly plausible in settings where programs are in close proximity to one another. It assumes that each member of a couple has their own preferences over individual programs, and their joint preferences are such that, when the assignment of one member is fixed, an improvement for the other member based on their individual preferences also improves the couple’s joint preferences. [^21] Responsive preferences were originally introduced by Roth (1985) in a two-sided matching market to derive institution preferences over groups of individuals from their preferences over individual members, in a class of settings without complementarities. In such settings, the individual-proposing DA algorithm, which disregards joint preferences and treats couples as separate individuals, generates a stable outcome.
@@
Subsequent literature illuminated the reasons behind the virtually guaranteed existence of stable outcomes in the computational experiments with simulated data reported in Roth and Peranson (1999). These experiments reflected the small proportion of residents in the NRMP pool who are members of dual-physician couples. Kojima et al. (2013) established that, with random preferences and a vanishingly small proportion of couples, a stable outcome almost always exists. However, in the same model, Ashlagi et al. (2014) showed that as the proportion of couples increases, the probability of a stable outcome existing sharply decreases.
@@
These more recent theoretical advances, inspired by practical challenges, illustrate how applied research can also contribute to fundamental understanding. This convergence of basic and applied research—use-inspired basic research—embodies the essence of Pasteur’s Quadrant. To further illustrate this, we turn to William Vickrey’s foundational work in auction theory.
3.2.3 Pasteur’s Quadrant: Vickey Auction and Revenue Equivalence Theorem